# Optimal Coupled Sensor Placement and Path-Planning in Unknown Time -Varying Environments

### v1

@@ 3 Coupled Sensing and Planning
Coupled sensor configuration and path-planning (CSCP) is our proposed iterative approach to solve Problem 1. At each iteration, a sensor configuration is determined, and measurements of the threat field are collected. The optimal sensor configuration is found by maximizing an information measure that we call context-relevant mutual information (CRMI). These sensor measurements are used to update the threat field estimates in an estimator. Next, the path plan is modified based on the new threat field estimate, and this process continues until the path cost variance is reduced below a prespecified threshold $\varepsilon$. In this paper we assume that the actor is a planning agent that does not move before the CSCP algorithm is complete. In other words, we may think of this implementation of CSCP as occurring in a virtual world for planning the future movements of the actor and sensors.
@@ 3.1 Context-Relevant Mutual Information (CRMI)
[…]
@@
$P_{{\boldsymbol{z}\boldsymbol{z}}_{k|k-1}}$ $=C_{k}(\boldsymbol{q})P_{\boldsymbol{\Theta}\boldsymbol{\Theta}_{k|k-1}}C_{k}^{\intercal}(\boldsymbol{q})+R_{k},$ (6)
$P_{{\boldsymbol{\Theta}\boldsymbol{z}}_{k|k-1}}$ $=P_{\boldsymbol{\Theta}\boldsymbol{\Theta}_{k|k-1}}C_{k}^{\intercal}(\boldsymbol{q}).$ (7)
The mutual information between the state $\boldsymbol{\Theta}_{k}$ and measurement $\boldsymbol{z}_{k}(\boldsymbol{q})$ is then written as [41]
@@
The MI $I(\boldsymbol{\Theta}_{k};\boldsymbol{z}_{k}(\boldsymbol{q}))$ depends on the sensor configuration $\boldsymbol{q},$ and in that sense it quantifies “informativeness” of a configuration $\boldsymbol{q}.$ A canonical method of finding the optimal sensor configuration is to maximize the MI $I$ over $\boldsymbol{q}.$ Note, however, that this MI $I(\boldsymbol{\Theta}_{k};\boldsymbol{z}_{k}(\boldsymbol{q}))$ is entirely decoupled from the path-planning problem, in that it is no way contextualized by the optimal path to be found. We introduce a new information measure called the context-relevant mutual information (CRMI), which couples the sensor configuration (placement) and path-planning problems.
We define CRMI as the mutual information between the path cost and the measurements. CRMI may be thought of as a formalization of the vague notion “sensors ‘near’ the planned path are more informative than those farther away.” To this end, consider that, for any path $\boldsymbol{v}$, the expected cost is $\widehat{J}(\boldsymbol{v}):=L+\delta\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\widehat{\boldsymbol{\Theta}}_{\ell}.$ The joint PDF $\textit{p}(J_{k},\boldsymbol{z}_{k})$ between the path cost and measurement variables is
$\textit{p}(J_{k},\boldsymbol{z}_{k})$ $=\mathcal{N}\left(\begin{bmatrix}J_{k}\\ \boldsymbol{z}_{k}\end{bmatrix}:\begin{bmatrix}\widehat{J}_{k|k-1}\\ \widehat{\boldsymbol{z}}_{k}\end{bmatrix},\begin{bmatrix}{P}_{JJ_{k|k-1}}&P_{{J\boldsymbol{z}}_{k|k-1}}\\ P^{\intercal}_{{J\boldsymbol{z}}_{k|k-1}}&P_{{\boldsymbol{z}\boldsymbol{z}}_{k|k-1}}\end{bmatrix}\right).$
@@
${P}_{JJ_{k|k-1}}$ $:=\mathbb{E}\left[\left(J(\boldsymbol{v})-\widehat{J}(\boldsymbol{v})\right)^{2}\right]=\mathbb{E}\left[\left(\delta\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\right)^{2}\right],$
which may be written as
@@
$P_{JJ_{k|k-1}}=\delta^{2}\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})^{\intercal}P_{k_{\ell}}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\right)+2\delta^{2}\sum_{\begin{subarray}{c}\ell<m\\ \ell,m\in[L]\end{subarray}}^{L}\left(\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})^{\intercal}P_{k_{\ell m}}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{m}})\right).$ (9)
The calculation of $P_{JJ_{k|k-1}}$ requires the determination of $\boldsymbol{\Phi}$ and the error covariance $P$ for every grid point $\boldsymbol{v}_{l}$ lying on the path. $P_{k_{\ell}}$ and $P_{k_{\ell m}}$ are determined by propagating the UKF prediction steps for time steps equivalent to the path length. The covariance of the measurement and the cross covariance between the path cost and the measurement random vector are formulated as,
@@
$P_{{J\boldsymbol{z}}_{k|k-1}}$ $=\mathbb{E}\left[\left(\boldsymbol{z}-\widehat{\boldsymbol{z}}\right)\left(J(\boldsymbol{v})-\widehat{J}(\boldsymbol{v})\right)\right]$ $=\delta\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})^{\intercal}P_{k_{\ell}}\right)C_{k}^{\intercal}(\boldsymbol{q}).$ (10)
@@
$P_{{\boldsymbol{z}\boldsymbol{z}}_{k|k-1}}$ $=\mathbb{E}\left[\left(\boldsymbol{z}-\widehat{\boldsymbol{z}}\right)\left(\boldsymbol{z}-\widehat{\boldsymbol{z}}\right)\right]$ $=C_{k}(\boldsymbol{q})P_{\boldsymbol{\Theta}\boldsymbol{\Theta}_{k|k-1}}C_{k}^{\intercal}(\boldsymbol{q})+R_{k}.$ (11)
@@ 3.2 CSCP Algorithm
The CSCP algorithm described in Algorithm 1 initializes with $\widehat{\boldsymbol{\Theta}}_{0}=\mathbf{0}$ and $P_{0}=\chi\boldsymbol{I}_{N_{\mathrm{P}}}$, where $\chi$ is a large arbitrary number. The initial sensor placement $\boldsymbol{q}_{0}$ is arbitrary. At the initial iteration, an optimal path $\boldsymbol{v}^{*}_{0}$ of minimum expected cost $\mathbb{E}\left[J(\boldsymbol{v}^{*}_{0})\right]$ is calculated.
@@
The optimal sensor configuration in Line 6 can be calculated by optimizing some measure of information gain. In a decoupled approach, we may optimize the standard MI in (8). In the proposed CSCP method, we optimize the CRMI in (12). The method of optimization is left to the user, and is not the focus of this paper. For a small number of grid points, we can determine $\boldsymbol{q}^{*}_{k}$ by mere enumeration.
@@
1:  Set $k=0,\widehat{\boldsymbol{\Theta}}_{0}=\mathbf{0},$ and $P_{0}=\chi\boldsymbol{I}_{N_{\mathrm{P}}}$
2:  Initialize sensor placement $\boldsymbol{q}_{0}\subset\{N_{\mathrm{G}}\}$
3:  Find $\boldsymbol{v}^{*}_{0}=\arg\min(\widehat{J}_{0}(\boldsymbol{v}))$
@@
5:     Determine $I({J}_{k};\boldsymbol{z}_{k}(\boldsymbol{q}))$ per (12)
@@
7:     Obtain new sensor measurements $\boldsymbol{z}_{k}(\boldsymbol{q}^{*}_{k})$
8:     Update $\widehat{\boldsymbol{\Theta}}_{k},P_{k}$
9:     Find $\boldsymbol{v}^{*}_{k}:=\arg\min(\widehat{J}_{k}(\boldsymbol{v}))$
10:     Increment iteration counter $k=k+1$
Figure 1: The proposed CSCP algorithm.

### v2

@@ 3 Coupled Sensing and Planning
Coupled sensor configuration and path-planning (CSCP) is our proposed iterative approach to solve Problem 1. At each iteration, a sensor configuration is determined, and measurements of the threat field are collected. The optimal sensor configuration is found by maximizing an information measure that we call context-relevant mutual information (CRMI). These sensor measurements are used to update the threat field estimates in an estimator. For future compatibility with nonlinear threat dynamics, an unscented Kalman filter (UKF) is used for the estimating the parameters $\boldsymbol{\Theta}.$ The reader interested is referred to Appendix B for a brief description of the UKF. Next, the path plan is modified based on the new threat field estimate, and this process continues until the path cost variance is reduced below a prespecified threshold $\varepsilon$.
In this paper, we assume that the actor is a planning agent that does not move before the CSCP algorithm is complete. In other words, we may think of this implementation of CSCP as occurring in a virtual world for planning the future movements of the actor and sensors. The reason for this restriction in scope is to be able to provide a proof of convergence of the CSCP method as a baseline for future applications. In our recent work [47], we implement a version of CSCP where the actor moves simultaneously with the CSCP iterations, and the sensors need a finite non-zero duration to relocate. The simulations in [47] demonstrate a successful implementation and provide sensor resource benefits similar to those described in the present manuscript, pending a formal proof of convergence.
@@ 3.1 Context-Relevant Mutual Information (CRMI)
$I(\boldsymbol{\Theta}_{k};\boldsymbol{z}_{k}):=\int\int\textit{p}(\boldsymbol{\Theta}_{k},\boldsymbol{z}_{k})\log\left(\frac{\textit{p}(\boldsymbol{\Theta}_{k},\boldsymbol{z}_{k})}{\textit{p}(\boldsymbol{\Theta}_{k})\textit{p}(\boldsymbol{z}_{k})}\right)\,d\boldsymbol{\Theta}_{k}\,d\boldsymbol{z}_{k},$
@@
$P_{{\boldsymbol{z}\boldsymbol{z}}_{k|k-1}}$ $:=\mathbb{E}\left[\left(\boldsymbol{z}_{k}-\widehat{\boldsymbol{z}}_{k}\right)\left(\boldsymbol{z}_{k}-\widehat{\boldsymbol{z}}_{k}\right)^{\intercal}\right]$
$=\mathbb{E}\left[\left(C_{k}(\boldsymbol{q})\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)+\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)\right)\left(C_{k}(\boldsymbol{q})\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)+\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)\right)^{\intercal}\right]$
$=C_{k}(\boldsymbol{q})\mathbb{E}\left[\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)^{\intercal}\right]C_{k}^{\intercal}(\boldsymbol{q})+\mathbb{E}\left[\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)^{\intercal}\right]$
$=C_{k}(\boldsymbol{q})P_{\boldsymbol{\Theta}\boldsymbol{\Theta}_{k|k-1}}C_{k}^{\intercal}(\boldsymbol{q})+R_{k},$ (6)
$P_{{\boldsymbol{\Theta}\boldsymbol{z}}_{k|k-1}}$ $:=\mathbb{E}\left[\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)\left(\boldsymbol{z}_{k}-\widehat{\boldsymbol{z}}_{k}\right)^{\intercal}\right]$
$:=\mathbb{E}\left[\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)\left(C_{k}(\boldsymbol{q})\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)+\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)\right)^{\intercal}\right]$
$=\mathbb{E}\left[\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)^{\intercal}\right]C_{k}^{\intercal}(\boldsymbol{q})+\mathbb{E}\left[\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k|k-1}\right)\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)^{\intercal}\right]$
$=P_{\boldsymbol{\Theta}\boldsymbol{\Theta}_{k|k-1}}C_{k}^{\intercal}(\boldsymbol{q}).$ (7)
The mutual information between the state $\boldsymbol{\Theta}_{k}$ and measurement $\boldsymbol{z}_{k}(\boldsymbol{q})$ is then written as [45]:
@@
The MI $I(\boldsymbol{\Theta}_{k};\boldsymbol{z}_{k}(\boldsymbol{q}))$ depends on the sensor configuration $\boldsymbol{q},$ and in that sense it quantifies “informativeness” of a configuration $\boldsymbol{q}.$ A canonical method of finding the optimal sensor configuration is to maximize the MI ($I$) over $\boldsymbol{q}.$ Note, however, that this MI $I(\boldsymbol{\Theta}_{k};\boldsymbol{z}_{k}(\boldsymbol{q}))$ is entirely decoupled from the path-planning problem, in that it is no way contextualized by the optimal path to be found. We introduce a new information measure called the context-relevant mutual information (CRMI), which couples the sensor configuration (placement) and path-planning problems.
[…]
@@
${P}_{JJ_{k|k-1}}$ $:=\mathbb{E}\left[\left(J(\boldsymbol{v})-\widehat{J}(\boldsymbol{v})\right)^{2}\right]=\mathbb{E}\left[\left(\delta\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\right)^{2}\right].$
Using the formula for square of sums, $\left(\sum_{i}a_{i}\right)^{2}=\sum_{i}a_{i}^{2}+2\sum_{i<j}a_{i}a_{j}$
@@
$P_{JJ_{k|k-1}}$ $=\delta^{2}\left(\mathbb{E}\left[\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\right)^{2}+2\sum_{\begin{subarray}{c}\ell<m\\ \ell,m\in\{L\}\end{subarray}}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{m}})\left(\boldsymbol{\Theta}_{m}-\widehat{\boldsymbol{\Theta}}_{m}\right)\right)\right]\right)$
@@
$=\delta^{2}\left(\mathbb{E}\left[\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\right)^{2}+2\sum_{\begin{subarray}{c}\ell<m\\ \ell,m\in\{L\}\end{subarray}}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\left(\boldsymbol{\Theta}_{m}-\widehat{\boldsymbol{\Theta}}_{m}\right)^{\intercal}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{m}})\right)\right]\right).$
@@
The first term $\mathbb{E}\left[\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\right)^{2}\right]$ can be further expressed as:
$P_{JJ_{k|k-1}}$ $=\sum_{\ell=1}^{L}\mathbb{E}\left[\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)^{\intercal}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\right]$
$=\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\mathbb{E}\left[\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)^{\intercal}\right]\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})=\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})P_{k_{\ell}}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}}).$
Therefore,
$P_{JJ_{k|k-1}}$ $=\delta^{2}\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})P_{k_{\ell}}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\right)+2\delta^{2}\sum_{\begin{subarray}{c}\ell<m\\ \ell,m\in\{L\}\end{subarray}}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})P_{k_{\ell m}}\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{v}_{m}})\right).$ (9)
The calculation of $P_{JJ_{k|k-1}}$ requires the determination of $\boldsymbol{\Phi}$ and the error covariance $P$ for every grid point $\boldsymbol{v}_{l}$ lying on the path. $P_{k_{\ell}}$ and $P_{k_{\ell m}}$ are determined by propagating the UKF prediction steps for iterations equivalent to the path length. The covariance of the measurement and the cross covariance between the path cost and the measurement random vector are formulated as,
$P_{{J\boldsymbol{z}}_{k|k-1}}$ $=\mathbb{E}\left[\left(J(\boldsymbol{v})-\widehat{J}(\boldsymbol{v})\right)\left(\boldsymbol{z}_{k}-\widehat{\boldsymbol{z}}_{k}\right)^{\intercal}\right]$
$=\mathbb{E}\left[\left(\delta\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\right)\left(C_{k}(\boldsymbol{q})\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k}\right)+\left(\boldsymbol{\eta}_{k}-\widehat{\boldsymbol{\eta}}_{k}\right)\right)^{\intercal}\right]$
$=\delta\mathbb{E}\left[\left(\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\left(\boldsymbol{\Theta}_{\ell}-\widehat{\boldsymbol{\Theta}}_{\ell}\right)\left(\boldsymbol{\Theta}_{k}-\widehat{\boldsymbol{\Theta}}_{k}\right)^{\intercal}C_{k}^{\intercal}(\boldsymbol{q})\right)\right]=\delta\sum_{\ell=1}^{L}\left(\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})P_{k_{\ell}}\right)C_{k}^{\intercal}(\boldsymbol{q}).$ (10)
$P_{{\boldsymbol{z}\boldsymbol{z}}_{k|k-1}}$ $=\mathbb{E}\left[\left(\boldsymbol{z}_{k}-\widehat{\boldsymbol{z}_{k}}\right)\left(\boldsymbol{z}_{k}-\widehat{\boldsymbol{z}_{k}}\right)^{\intercal}\right]=C_{k}(\boldsymbol{q})P_{\boldsymbol{\Theta}\boldsymbol{\Theta}_{k|k-1}}C_{k}^{\intercal}(\boldsymbol{q})+R_{k}.$ (11)
@@ 3.2 CSCP Algorithm
The CSCP algorithm described in Algorithm 1 initializes with $\widehat{\boldsymbol{\Theta}}_{0}=\mathbf{0}$ and $P_{0}=\chi\boldsymbol{I}_{N_{\mathrm{P}}}$, where $\chi$ is a large arbitrary number. The initial sensor placement $\boldsymbol{q}_{0}$ is arbitrary. At the initial iteration, an optimal path $\boldsymbol{v}^{*}_{0}$ of minimum expected cost $\mathbb{E}\left[J(\boldsymbol{v}^{*}_{0})\right]$ is calculated. This threat estimate initialization implies that until a measurement is taken, the algorithm assumes all threat states to be zero. As a result, the CSCP planner is “optimistic” in that it plans paths through regions associated with threat states that were not previously measured or estimated. The CRMI-based sensor placement then ensures sensors are placed close to this path, thereby ensuring that the sensors explore the entire workspace.
@@
The optimal sensor configuration in Line 6 can be calculated by optimizing some measure of information gain. In a decoupled approach, we may optimize the standard MI in (8). In the proposed CSCP method, we optimize the CRMI in (12). The method of optimization is left to the user, and is not the focus of this paper. For a small number of grid points, we can determine $\boldsymbol{q}^{*}_{k}$ by mere enumeration. In prior works, e.g., [6], we have found success in implementing evolutionary global optimization methods for sensor configuration, albeit using a different reward function. It is possible to apply such methods for CRMI maximization as well.
@@
1: Set $k=0,\widehat{\boldsymbol{\Theta}}_{0}=\mathbf{0},$ and $P_{0}=\chi\boldsymbol{I}_{N_{\mathrm{P}}}$;
2: Initialize sensor placement $\boldsymbol{q}_{0}\subset\{N_{\mathrm{G}}\}$;
3: Find $\boldsymbol{v}^{*}_{0}=\arg\min(\widehat{J}_{0}(\boldsymbol{v}))$;
@@
5:  Determine $I({J}_{k};\boldsymbol{z}_{k}(\boldsymbol{q}))$ per (12);
@@
7:  Obtain new sensor measurements $\boldsymbol{z}_{k}(\boldsymbol{q}^{*}_{k})$;
8:  Update $\widehat{\boldsymbol{\Theta}}_{k},P_{k}$ ;
9:  Find $\boldsymbol{v}^{*}_{k}:=\arg\min(\widehat{J}_{k}(\boldsymbol{v}))$;
10:  Increment iteration counter $k=k+1$.
Algorithm 1 Coupled Sensor Configuration and Planning (CSCP)